# Auto-Search and Refinement: An Automated Framework for Gender Bias Mitigation in Large Language Models

### v2

@@ 1 Introduction
Existing gender debiasing methods struggle to fulfill all these requirements simultaneously. Efforts to align LLMs with bias-free values include parameter-modification methods such as supervised fine-tuning, reinforcement learning on human preference data [38, 32, 49, 1], or model editing on specific examples [6, 2]. However, these approaches face limitations in accessibility, efficiency, and flexibility, making them unsuitable for API-based models and requiring significant computational resources as models scale. Instruction-based approaches provide an alternative solution, leveraging the instruction-following capability of LLMs, where preambles are appended to counteract gender stereotypes or reframe concepts in a gender-neutral manner to reduce bias. Debiasing prompts can be manually designed using counterfactual, fairness requirements and descriptive preambles [29, 48], or automatically generated through gradient-based methods that expand the prompt search space [35, 34]. However, automated searching for debiasing prompts requires white-box access to the model, while manually designed prompts often lack automation and compromise the LLMs’ performance on normal tasks. For example, Oba et al. [27] introduces appending preambles before the input text to mitigate gender bias. These preambles include detailed counterfactual statements (CF-D), which counteract real-world stereotypical gender associations, and detailed descriptive (Desc-D) preambles, which reframe gender-biased concepts in a gender-neutral manner. While effective in reducing bias, these preambles may unintentionally degrade the model’s performance on normal tasks. As observed in Figure 1, both open-source and closed-source LLMs show reduced response quality on standard benchmarks like Dolly. This decline stems from the preamble’s distorting the model’s interpretation of the query. For instance, as shown in Table 1, the manually designed counterfactual statement CF-D introduces unnecessary gender-related details, which could mislead the model on gender-irrelevant queries.
@@
Specifically, as illustrated in Figure 2, the auto-search phase begins by optimizing bias-reducing triggers using a preference dataset, then filtering based on their debiasing effectiveness to construct a Fairwords Bag and a corresponding preference dataset. In the refinement phase, a sequence-to-sequence (seq2seq) model is trained to adaptively refine Fairwords for diverse input queries from both normal tasks and gender-related tasks, enhancing FaIRMaker’s transferability and allowing it to function as an independent module. The refiner is designed to generate context-aware Fairwords: it preserves model performance when handling neutral queries, and produces effective debiasing instructions when facing bias-related queries. To balance bias mitigation and normal task performance, the specialized refiner is trained on the LLM-assisted refined-Fairwords dataset that contains queries from both normal tasks and gender-related tasks. During inference, FaIRMaker randomly selects a Fairwords from the Fairwords Bag and uses the refiner to generate the refined Fairwords as an instruction to guide the model’s response based on the input query. This process ensures fairness without degrading performance on neutral queries. Experimental results demonstrate that FaIRMaker outperforms baseline methods on both open- and closed-sourced LLMs in terms of gender bias mitigating effectiveness and utility on normal tasks.
@@
  • Extensive experiments on closed- and open-source LLMs, such as GPT series, Qwen series, and llama2 series demonstrate the effectiveness of FaIRMaker on mitigating bias while preserving task integrity across diverse downstream tasks, as well as its efficiency, extendability and interpretability analysis. Comprehensive ablation studies reveal each component’s contribution.
@@ 2.1 Gender Bias in LLMs
Gender bias in LLMs can be evaluated through intrinsic and extrinsic approaches [20, 46]. Intrinsic methods evaluate bias independent of specific downstream tasks by analyzing statistical associations in the embedding space [18, 24] or evaluating the probabilities assigned to different options in datasets [26, 25]. In contrast, extrinsic approaches examine gender bias within the context of downstream tasks, such as coreference resolution [19, 17], question answering [13], reference letter generation [40], and classification tasks [10], each capturing gender bias from distinct perspectives. These studies underscore the need for ongoing research and mitigation strategies.
@@ 2.2 Gender Bias Mitigation in LLMs
To address gender bias in LLMs, various strategies have been proposed, typically categorized into white-box and black-box methods based on access to a model’s internal parameters. White-box methods require access to internal parameters, including fine-tuning and model editing. Fine-tuning involves creating specialized gender-inclusive datasets [3, 11] for instruction-based fine-tuning [32, 38] or Direct Preference Optimization (DPO; 49, 1). Model editing focuses on identifying and modifying bias pathways [6] or utilizing hyper-networks for automatic parameter updates [2]. While effective, these methods depend on parameter access, limiting their use to closed-source models and potentially impacting overall model performance.
Black-box methods mitigate bias without requiring parameter access, often using textual prompts to guide fairer outputs. Techniques such as Chain of Thought (CoT; 43) and in-context learning (ICL; 5) have shown considerable promise [33, 14]. Counterfactual prompts and curated examples effectively encourage equitable content generation [36, 12, 27]. However, they rely on static prompts, which may lose effectiveness on novel tasks or out-of-distribution data, limiting their robustness.
@@ 3.1 Fairwords Auto-Searching
Fairwords filtering. The Fairwords filtering process evaluates whether the Fairwords identified in the optimization step genuinely reduce gender bias. Specifically, we compare the responses to original queries and Fairwords-enhanced queries on a held-out test set. The llama3.1-8b-instruct model serves as a judge, assessing both response quality and bias levels using a predefined evaluation prompt (see Appendix F.2). Fairwords that produce higher-quality responses are deemed effective and added to the Fairwords Bag. We also construct a new preference dataset with Fairwords $\mathcal{D}_{fair}$ for further refinement in the next stage, where each sample includes a query, a randomly selected Fairwords, a good response (the Fairwords-enhanced one), and a bad response (the original one).
@@ 3.2 Instruction Generator Training
Although the filtered Fairwords can prompt better-quality responses, they are nonsensical token combinations lacking interpretability and transferability across black-box LLMs. Additionally, model performance on standard tasks should be maintained.
To ensure FaIRMaker is a model-independent module compatible with both open-source and API-based LLMs while preserving their original performance on normal tasks, we introduce a refinement step. This step transforms the unintelligible Fairwords into human-readable prompts, performing a reverse inference process on the preference dataset of both tasks with the assistance of ChatGPT. Then, a seq2seq model is trained to generalize and learn how to refine the Fariword to mitigate bias and execute normal tasks without the preference dataset. As a result, given any query and Fairwords, FaIRMaker adaptively generates refined Fairwords, ensuring robust performance on both bias mitigation and task execution without compromising utility.
@@ 4.1 Configurations
Evaluation Metrics. To evaluate gender bias mitigation on GA-test, we use win-tie-loss rates. Following prior work [41, 51], DeepSeek-V3 (DS) [22], Gemini-2.0-Flash (Gemini) [15], and GPT4 act as a judge to score responses based on a predefined evaluation prompt (see Appendix F.2). We compare the scores of bias-mitigated and original responses, reporting win, tie, and lose proportions. For BBQ-gender, we adopt the sDIS and sAMB metrics to measure the gender bias in disambiguated and ambiguous contexts respectively, which is defined in the original paper (See Appendix C for detailed definition). For MMLU and HellaSwag, we use the accuracy as the evaluation metric. In utility datasets involving open-ended QA tasks, response score (RS) judged by evaluators is used for performance evaluation with a custom prompt (Appendix F.2). We also measure the time cost per query to assess efficiency. All the results are averaged over four random seeds.
@@ 4.2 Bias Mitigation
Figure 3: Performance comparison between the base models and the models after applying FaIRMaker on the GA-test dataset.
@@ 4.5 Extendability
FaIRMaker functions as an independent inference-time module, enabling integration with other debiasing methods such as applying Direct Preference Optimization (DPO; 31) on bias-free datasets. This section demonstrate the extendability of FaIRMaker through comparing the performance of FaIRMaker and DPO, and explores their potential synergy when combined.
FaIRMaker v.s. DPO. We fine-tune the Llama2-Alpaca model on the GenderAlign (GA) dataset using DPO [49]. As shown in Figure 6, DPO-based method demonstrates better performance on in-distribution data (GA-test) and struggles on out-of-distribution generalization, performing worse on BBQ-gender compared to FaIRMaker. Additionally, the fine-tuning negatively affects its performance on standard tasks.
@@ 4.6 Ablation Study
In this section, we evaluate the contributions of each FaIRMaker module through ablation studies. We define three variants: (1) w/o auto-search, which skips the automatic FairWords discovery step and instead uses ChatGPT to directly generate refinement instructions from preference data, (2) w/o filtering, where all Fairwords and responses are used without filtering in auto-search, (3) w/o refinement, where Fairwords from the Fairwords Bag are directly appended to queries without refinement. We evaluate these ablations on Llama2-Alpaca for bias mitigation and general tasks. Table 6 presents the results, with the best highlighted in bold.
@@
win rate (GA-test) ($\uparrow$) 53.15% 50.64% 51.94% 42.79%
sDIS (BBQ-gender) ($\downarrow$) 0.224 0.189 0.561 0.675
@@
Role of Auto-Searching: While directly generating and using prompts from ChatGPT can mitigate bias to a certain extent, this approach generally underperforms compared to FaIRMaker. These results suggest that the Fairwords produced during the auto-searching phase play a crucial role by serving as effective seed signals that guide the refinement process more reliably and consistently. Role of Filtering: The filtering step in the auto-search ensures that only Fairwords with genuine debiasing effects move to the next stage. FaIRMaker w/o filtering shows reduced bias mitigation and inconsistent performance on general tasks. Without filtering, noisy gender-related data disrupts the refiner’s training, impairing feature extraction and weakening bias mitigation effectiveness. Role of Refinement: The refinement step converts Fairwords into natural language instructions, enhancing FaIRMaker’s generalization and transferability to black-box models. FaIRMaker w/o refinement exhibits significantly lower performance, indicating its limitations in generalization.
@@ 4.7 Interpretation
The vanilla Fairwords optimized in the auto-search step are often nonsensical token combinations aimed at maximizing favorable and minimizing unfavorable responses. Their role as debiasing triggers or task enhancers remains unclear. To explore this, we prompted both ChatGPT and Llama2-Alpaca to interpret their potential meanings and emotional connotations (see Table 7). ChatGPT often associated Fairwords with emotions like urgency and seriousness, while Llama2-Alpaca responses were more expressive, frequently conveying frustration (54.84%) and anger (31.18%). These emotional cues may steer models toward less biased responses while preserving the input’s original intent due to the semantic ambiguity of Fairwords.
@@ Appendix B Detailed Experiment Configuration
Fairwords searching setup. In our experiment, we set the length of the Fairwords $l$ to 20, the batch size $m$ to 25, the number of top-weight candidates $b$ to 256, and the sampling size $k$ to 512. All searches are performed on a single NVIDIA A40 GPU, with the optimization process taking around 24 hours to complete 300 steps.

### v3

@@ 1 Introduction
Existing gender debiasing methods struggle to fulfill all these requirements simultaneously. Efforts to align LLMs with bias-free values include parameter-modification methods such as supervised fine-tuning, reinforcement learning on human preference data [44, 37, 55, 1], or model editing on specific examples [7, 2]. However, these approaches face limitations in accessibility, efficiency, and flexibility, making them unsuitable for API-based models and requiring significant computational resources as models scale. Instruction-based approaches provide an alternative solution, leveraging the instruction-following capability of LLMs, where preambles are appended to counteract gender stereotypes or reframe concepts in a gender-neutral manner to reduce bias. Debiasing prompts can be manually designed using counterfactual, fairness requirements, and descriptive preambles [34, 54], or automatically generated through gradient-based methods that expand the prompt search space [41, 40]. However, automated searching for debiasing prompts requires white-box access to the model, while manually designed prompts often lack automation and compromise the LLMs’ performance on normal tasks. For example, Oba et al. [32] introduces appending preambles before the input text to mitigate gender bias. These preambles include detailed counterfactual statements (CF-D), which counteract real-world stereotypical gender associations, and detailed descriptive (Desc-D) preambles, which reframe gender-biased concepts in a gender-neutral manner. While effective in reducing bias, these preambles may unintentionally degrade the model’s performance on normal tasks. As observed in Figure 1, both open-source and closed-source LLMs show reduced response quality on standard benchmarks like Dolly. This decline stems from the preamble’s distorting the model’s interpretation of the query. For instance, Table 1 illustrates that the manually designed counterfactual statement CF-D introduces extraneous gender cues, leading the model to produce biased responses to gender-neutral queries.
@@
Specifically, as illustrated in Figure 2, the auto-search phase begins by optimizing bias-reducing triggers using a preference dataset, then filtering based on their debiasing effectiveness to construct a Fairwords Bag and a corresponding preference dataset. In the refinement phase, a sequence-to-sequence (seq2seq) model is trained to adaptively refine Fairwords for diverse input queries from both normal tasks and gender-related tasks, enhancing FaIRMaker’s transferability and allowing it to function as an independent module. The refiner is designed to generate context-aware Fairwords: it preserves model performance when handling neutral queries, and produces effective debiasing instructions when facing bias-related queries. To balance bias mitigation and normal task performance, the specialized refiner is trained on the LLM-assisted refined-Fairwords dataset that contains queries from both normal tasks and gender-related tasks. During inference, FaIRMaker randomly selects a Fairwords from the Fairwords Bag and uses the refiner to generate the refined Fairwords as an instruction to guide the model’s response based on the input query. This process ensures fairness without degrading performance on neutral queries.
@@
  • Extensive experiments on closed- and open-source LLMs, such as GPT series, Qwen series, and llama2 series demonstrate the effectiveness of FaIRMaker on mitigating bias while preserving task integrity across diverse downstream tasks, as well as its efficiency, extendability and interpretability analysis. Comprehensive ablation studies reveal each component’s contribution [^1] The code is available at https://github.com/SavannahXu79/FaIRMaker.
@@ 2.1 Gender Bias in LLMs
Gender bias in LLMs can be evaluated through intrinsic and extrinsic approaches [23, 52]. Intrinsic methods evaluate bias independent of specific downstream tasks by analyzing statistical associations in the embedding space [21, 28] or evaluating the probabilities assigned to different options in datasets [31, 30]. In contrast, extrinsic approaches examine gender bias within the context of downstream tasks, such as coreference resolution [22, 20], question answering [15], reference letter generation [46], and classification tasks [12], each capturing gender bias from distinct perspectives. These studies underscore needs for ongoing research and mitigation strategies.
@@ 2.2 Gender Bias Mitigation in LLMs
To address gender bias in LLMs, various strategies have been proposed, typically categorized into white-box and black-box methods based on access to a model’s internal parameters. White-box methods require access to internal parameters, including fine-tuning and model editing. Fine-tuning involves creating specialized gender-inclusive datasets [4, 13] for instruction-based fine-tuning [37, 44] or Direct Preference Optimization (DPO; 55, 1). Model editing focuses on identifying and modifying bias pathways [7] or utilizing hyper-networks for automatic parameter updates [2]. While effective, these methods depend on parameter access, limiting their use to closed-source models and potentially impacting overall model performance.
Black-box methods mitigate bias without requiring parameter access, often using textual prompts to guide fairer outputs. Techniques such as Chain of Thought (CoT; 49) and in-context learning (ICL; 6) have shown considerable promise [38, 16]. Counterfactual prompts and curated examples effectively encourage equitable content generation [42, 14, 32]. However, they rely on static prompts, which may lose effectiveness on novel tasks or out-of-distribution data, limiting their robustness.
@@ 3.1 Fairwords Auto-Searching
Fairwords filtering. The Fairwords filtering process evaluates whether the Fairwords identified in the optimization step genuinely reduce gender bias. Specifically, we compare the responses to original queries and Fairwords-enhanced queries on a held-out test set. The llama3.1-8b-instruct model is employed as the evaluator owing to its comparatively low gender bias reported in prior studies [3, 19] and its lightweight nature, which enables scalable evaluation. It assesses both response quality and bias levels using a predefined evaluation prompt (see Appendix F.2). To further reduce noise, each response is evaluated three times, and only pairs with a score margin greater than 0.5 are retained. Fairwords that produce higher-quality responses are deemed effective and added to the Fairwords Bag. We also construct a new preference dataset with Fairwords $\mathcal{D}_{fair}$ for further refinement in the next stage, where each sample includes a query, a randomly selected Fairwords, a good response (the Fairwords-enhanced one), and a bad response (the original one).
@@ 3.2 Instruction Generator Training
Although the filtered Fairwords can elicit higher-quality responses, they are often nonsensical token sequences that lack interpretability and transferability across black-box LLMs [9]. Moreover, it is essential to maintain model performance on standard tasks. To make FaIRMaker a model-agnostic module compatible with both open-source and API-based LLMs while preserving their original task performance, we introduce a refinement step. This step transforms unintelligible Fairwords into human-readable prompts through a reverse-inference process conducted on the preference datasets of both tasks, assisted by ChatGPT. Such semantic transformation is a common practice in prompt engineering [8, 29] and has been shown to preserve the effectiveness of optimization-based suffixes [25]. We hypothesize that LLMs can internalize and retain the behavioral intent of these raw Fairwords even after natural-language rewriting. Subsequently, a sequence-to-sequence model is trained to generalize this refinement process, learning to convert Fairwords into readable prompts that mitigate bias and perform standard tasks without access to preference datasets. Consequently, given any query and Fairwords, FaIRMaker adaptively generates refined instructions that ensure robust performance in both bias mitigation and task execution without compromising overall utility.
@@ 4.1 Configurations
Evaluation Metrics. To evaluate gender bias mitigation on GA-test, we use win-tie-loss rates. Following prior work [47, 57], DeepSeek-V3 (DS) [26], Gemini-2.0-Flash (Gemini) [17], and GPT4 act as a judge to score responses based on a predefined evaluation prompt (see Appendix F.2). We compare the scores of bias-mitigated and original responses, reporting win, tie, and loss proportions. For BBQ-gender, we adopt the sDIS and sAMB metrics to measure the gender bias in disambiguated and ambiguous contexts respectively, which are defined in the original paper (See Appendix C for detailed definition). For MMLU and HellaSwag, we use the accuracy as the evaluation metric. In utility datasets involving open-ended QA tasks, response score (RS) judged by evaluators is used for performance evaluation with a custom prompt (Appendix F.2). We also measure the time cost per query to assess efficiency. All the results are averaged over four random seeds.
@@ 4.2 Bias Mitigation
Figure 3: Performance comparison between the base models and the models after applying FaIRMaker on the GA-test dataset, evaluated by DeepSeek-V3, Gemini-2.0-Flash and GPT-4.
@@ 4.5 Extendability
FaIRMaker functions as an independent inference-time module, enabling integration with other debiasing methods such as applying Direct Preference Optimization (DPO; 36) on bias-free datasets. This section demonstrates the extendability of FaIRMaker through comparing the performance of FaIRMaker and DPO, and explores their potential synergy when combined.
@@
FaIRMaker v.s. DPO. We fine-tune the Llama2-Alpaca model on the GenderAlign (GA) dataset using DPO [55]. As shown in Figure 6, the DPO-based method demonstrates better performance on in-distribution data (GA-test) and struggles on out-of-distribution generalization, performing worse on BBQ-gender compared to FaIRMaker. Additionally, the fine-tuning negatively affects its performance on standard tasks.
@@ 4.6 Ablation Study
In this section, we conduct a series of ablation studies to analyze the contributions of each FaIRMaker module and the diversity of Fairwords in the Fairwords Bag.
Component Analysis. To evaluate the contributions of each FaIRMaker module, we define three variants: (1) w/o auto-search, which skips the automatic FairWords discovery step and instead uses ChatGPT to directly generate refinement instructions from preference data, (2) w/o filtering, where all Fairwords and responses are used without filtering in auto-search, (3) w/o refinement, where Fairwords from the Fairwords Bag are directly appended to queries without refinement. We evaluate these ablations on Llama2-Alpaca for bias mitigation and general tasks. Table 6 presents the results, with the best highlighted in bold.
@@
win rate (GA-test) ($\uparrow$) 54.23% 50.64% 51.94% 42.79%
sDIS (BBQ-gender) ($\downarrow$) 0.024 0.189 0.561 0.675
@@
Role of Auto-Searching: While directly generating and using prompts from ChatGPT can mitigate bias to a certain extent, this approach generally underperforms compared to FaIRMaker. These results suggest that the Fairwords produced during the auto-searching phase play a crucial role by serving as effective seed signals that guide the refinement process more reliably and consistently. Role of Filtering: The filtering step in the auto-search ensures that only Fairwords with genuine debiasing effects move to the next stage. FaIRMaker w/o filtering shows reduced bias mitigation and inconsistent performance on general tasks. Without filtering, noisy gender-related data disrupts the refiner’s training, impairing feature extraction and reducing bias mitigation effectiveness as well. Role of Refinement: The refinement step converts Fairwords into natural language instructions, enhancing FaIRMaker’s generalization and transferability to black-box models. FaIRMaker w/o refinement exhibits significantly lower performance, indicating its limitations in generalization. Additionally, we evaluate the log-probability gap ($\Delta p=p_{chosen}-p_{rejected}$) between favorable and unfavorable responses under the guidance of refined versus vanilla Fairwords. As shown in Table 7, the results for refined Fairwords remain positive on the target model Llama2-Alpaca and consistently surpass those of vanilla Fairwords across other transfer models.
Avg. $\Delta p$ (std.) Llama2-Alpaca Qwen2-Instruct Qwen2.5-Instruct GPT3.5-turbo
vanilla Fairwords ($\uparrow$) 5.42 (15.27) 6.02 (48.25) 5.91 (42.59) 8.23 (58.85)
refined Fairwords ($\uparrow$) 1.42 (34.65) 6.42 (55.07) 6.15 (52.67) 10.00 (67.08)
Table 7: Log-probability gap between good and bad responses under the guidance of refined versus vanilla Fairwords.
Fairwords Diversity. Our auto-searched Fairwords set consists of 93 items, which are consistently used across all experiments in this paper. We evaluate their semantic and lexical diversity using sentence embeddings, K-Means clustering, and BLEU scores. Specifically, pairwise cosine similarities between Fairwords tokens are computed using all-MiniLM-L6-v2222https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 embeddings, yielding an average similarity of 0.2691 with a standard deviation of 0.1157, which indicates substantial semantic spread rather than redundancy. To examine the clustering structure, we apply K-Means clustering to the embeddings and evaluate cluster quality using the Silhouette Score [39]. The results, summarized in Table 8, show a steady increase in Silhouette Score as $k$ varies from 2 to 10, peaking at $k=10$ with an average cluster size of approximately nine Fairwords. This suggests that the Fairwords set naturally forms at least ten semantically distinct groups, consistent with diverse mitigation strategies. We refrain from using higher $k$ values to avoid overfragmenting the Fairword space into overly fine clusters. Furthermore, the average pairwise BLEU score is 0.0068, confirming extremely low lexical overlap among Fairwords. Together, these findings demonstrate that our Fairwords set is both semantically and lexically diverse, supporting its generalizability and robustness across models.
$\bm{k}$ 2 3 4 5 6 7 8 9 10
Silhouette Score 0.0548 0.0463 0.0471 0.0500 0.0547 0.0561 0.0724 0.0811 0.0835
Table 8: Silhouette Scores of K-Means clustering on Fairwords embeddings.
@@ 4.7 Interpretation
To better understand the functional role of Fairwords, we analyze the semantics and emotional connotations underlying the automatically optimized token sequences. Although the vanilla Fairwords produced during the auto-search step effectively enhance debiasing performance, they are often nonsensical token combinations designed purely to maximize favorable responses while minimizing unfavorable ones. The mechanisms through which these Fairwords act as debiasing triggers or task enhancers therefore, remain unclear.
To gain insight, we prompt both ChatGPT and Llama2-Alpaca to interpret the potential meanings and emotional tones of several representative Fairwords (see Table 9). Interestingly, ChatGPT often associates Fairwords with emotions such as urgency and seriousness, whereas Llama2-Alpaca produced more expressive interpretations, frequently reflecting frustration (54.84%) and anger (31.18%). These emotional cues may implicitly steer models toward less biased yet intent-aligned responses, as the semantic ambiguity of Fairwords allows flexible modulation of model behavior without altering task semantics.
@@ 5 Conclusion
Acknowledgement
We sincerely thank the reviewers for their insightful and constructive comments. This work is supported by the Shanghai Engineering Research Center of Intelligent Vision and Imaging, the Open Research Fund of the State Key Laboratory of Blockchain and Data Security, Zhejiang University.
@@ Appendix B Detailed Experiment Configuration
Fairwords searching setup. In our experiment, we set the length of the Fairwords $l$ to 20, 30, and 40, the batch size $m$ to 25, the number of top-weight candidates $b$ to 256, and the sampling size $k$ to 512. All searches are performed on a single NVIDIA A40 GPU, with the optimization process taking around 24 hours to complete 300 steps.
@@ D.5 Case study
D.6 Human Evaluation
Dataset (Judge) Acc (%) FPR (%) FNR (%)
$\mathcal{D}_{fair}$ (Llama3.1) 86 8 6
GA-test (GPT-4) 90 4 6
Dolly Eval (GPT-4) 84 6 10
Table 14: Human-LLM evaluation alignment.
To verify that the LLM-as-a-judge evaluations align with human judgment, we conduct a small-scale human study comparing the decisions of GPT-4 and Llama3.1-8b-instruct, which serve as the primary evaluators in our experiments and filtering process. Specifically, 50 preference pairs are randomly selected, and the decisions made by four human annotators (two female and two male) are compared with those of the LLM judges. As shown in Table 14, the results show strong agreement between human and LLM evaluations in terms of Accuracy, False Positive Rate (FPR), and False Negative Rate (FNR), indicating that the LLM-based evaluation process is both consistent and reliable.